# Universal Hitchin moduli spaces

### v1

@@ 3.3. Minimal coupling and Hamiltonian action
For the second equality, we have used that, for any vector field $y$ on $\sigma$, one has $L_{y}J=2i\bar{\partial}y^{1,0}-2i\partial y^{0,1}$ and therefore
$B(J\psi\wedge\psi(L_{y}J))$ $=2B(\psi^{1,0}\wedge\psi(\bar{\partial}y^{1,0}))+cc.$
$=2B(\psi^{1,0}\wedge(i_{y^{1,0}}(\bar{\partial}_{A}\psi^{1,0})+\bar{\partial}_{A}(i_{y^{1,0}}\psi)))+cc.$
$=-2\bar{\partial}(B(\psi^{1,0},\psi(y^{1,0})))+4B(\bar{\partial}_{A}\psi^{1,0},\psi(y^{1,0}))+cc.$
The $\widetilde{{\curly K}}-$action produces a ‘coupling term’ in the base $\Sigma$ (cf. Remark 4.2), which combines with the scalar curvature of the metric $g=\omega(,J)$. In particular, similarly as in [AGG], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}$ are given by solutions of the coupled system of equations
(3.23) $\begin{split}d_{A}^{*}\psi&=0\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
where $\chi(\Sigma)$ is the Euler characteristic of $\Sigma$ and $V=\int_{\Sigma}\omega$ is the total volume. In the next section we combine these equations with the flatness condition $F_{D}=0$ for the complex connection, in order to introduce the universal Hitchin moduli space.
@@ 4.1. Universal moduli space of flat $\boldsymbol{G}$-connections
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Let $\curly D$} be the space of connection on $E_{G}$, equipped with the constant complex structure (2.1). Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the Kähler fibration $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ defined in Section 3.2. Our aim in this section is to investigate the existence of a Kähler structure on $\curly{X}$ and quotients thereof, combining the minimal coupling for the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6, as considered in Section 3.3, with symplectic reduction. Difficulties will arise from Corollary 3.7, which shows that the associated symmetric tensor $\mathbf{g}_{\mathbb{J}}:=\sigma(,\mathbb{J})$ is negative semi-definite along horizontal directions.
@@
In different words, the extension (4.1) is given by restriction of the natural map from holomorphic automorphisms of $\curly{X}$ which preserve the fibration structure, to holomorphic automorphisms of the base $\curly{J}$. It is important to observe that even though $\operatorname{Diff}_{0}(\Sigma)$ acts by holomorphic automorphisms of the base $\curly{J}$, this is a real infinite-dimensional Lie group, which does not admit a complexification (see e.g. [Do3]).
@@
Similarly as in the previous section, $\Omega_{\mathbb{J}}$ induces a structure of holomorphic symplectic fibration on $\curly{X}\to\curly{J}$ and a holomorphic Ehresmann connection, which in this case is simply the trivial connection. We will return to this structure when studying the universal Higgs bundle moduli space in Section 5. The following result is an immediate consequence of the proof of [AGG, Proposition 1.6], but we sketch the proof in the present setup for the convenience of the reader.
@@
where $D\zeta\in\Omega^{0}(X,E_{G}(\mathfrak{g}))$ denotes the vertical part of the $G$-equivariant vector field $\zeta$ on the total space of $E_{G}$, with respect to the connection $D$.
@@
On the other hand, the infinitesimal action of $\zeta$ on $D$ is given by
$\zeta\cdot D=-d_{D}(D\zeta)-i_{p\zeta}F_{D}$
and the result follows. ∎
It is remarkable that, unlike in the moment calculation in Proposition 3.9, in the present setup the extension of the complex gauge group by diffeomorphisms does not produce any additional ‘coupling term’ in the base $\Sigma$ (cf. Remark 4.2). One can intepret this fact as a functorial property of flat connections under the action of diffeomorphisms on a surface. This allows us to define the universal moduli space of flat $G$-connections, as follows, and to impose the flatness condition for $D$ as an integrability condition in the next sections.
@@
$\mu_{\widetilde{\curly{G}}|\curly{X}^{*}}\colon\curly{X}^{*}\to\operatorname{Lie}\widetilde{\curly{G}}^{*}$.
@@
By construction, $\mathcal{U}^{\operatorname{Flat}}(G)$ fibers over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
(4.2) $\mathcal{U}^{\operatorname{Flat}}(G)\to\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
@@
For a $G$-bundle on a higher dimensional symplectic manifold $(M^{2n},\omega)$, consider the subgroup $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$ given by the preimage by $p$ of the group of Hamiltonian symplectomorphisms on $M$. Then, the proof of [AGG, Proposition 1.6] shows that the induced $\widetilde{{\curly K}}-$action on $\curly D$} is Hamiltonian, with equivariant moment map
@@
where $f\in C^{\infty}(M,\mathbb{R})$ is the unique smooth function satisfying $i_{y}\omega=df$ and $\int_{M}f\omega^{n}=0$, for $p(\zeta)=y$. In this case, there is a ‘coupling term’ in the base corresponding to the function
@@ 4.2. Universal moduli space of the coupled harmonic equations
In this section we combine the results of the previous two sections in order to define universal moduli spaces of solutions of the harmonicity equations (2.14), varying over Teichmuller space. As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
@@
We say that a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ is a solution of the coupled harmonic equations with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, if the following conditions are satisfied
@@
As in Section 4.1, consider ${\curly D}^{*}\subset{\curly D}$ the complex subspace of reductive flat connections. Then, $\curly{X}^{*}:=\curly{J}\times{\curly D}^{*}$ is formally a complex submanifold of $\curly{X}$ preserved by the $\widetilde{{\curly K}}-$action (see Section 3.3), and inherits a minimal coupling structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ and moment map $\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}$. We define the universal Hitchin moduli space as the symplectic quotient
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}:=\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}^{-1}(0)/\widetilde{{\curly K}}$,
By construction, $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ parametrizes solutions of the coupled harmonic equations (4.3) with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, modulo the $\widetilde{{\curly K}}-$action, and inherits a natural pre-symplectic structure on its smooth locus. This structure is furthermore symplectic in the case $\varepsilon=-1$, by Lemma 3.8. There are natural maps (cf. (5.3))
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)\to\mathcal{T}.$
@@
(4.4) $(\dot{J},a,\dot{\psi})\in\mathcal{S}^{1}:=T_{J}\curly{J}\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(X,E_{K}(\mathfrak{k})).$
@@
We denote by ${\bf L}_{\alpha}^{\varepsilon}(\dot{J},a,\dot{\psi})$ the differential operator defined by the left-hand side of equations (5.17). We turn next to the study of the infinitesimal action, in order to define a complex. From the proof of Proposition 3.9, using the connection $A$ we can identify elements $\zeta\in\operatorname{Lie}\widetilde{{\curly K}}$ with pairs
$\zeta\cong(f,u)\in\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k})),$
@@
$\mathcal{S}^{2}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus(\Omega^{2}(X,E_{K}(\mathfrak{k}))^{\oplus 3},$
@@
(4.7) ${(\mathcal{S}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{S}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}}$${\mathcal{S}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L}_{\alpha}^{\varepsilon}}$${\mathcal{S}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
The cohomology $H^{1}(\mathcal{S}^{*}):=\frac{\operatorname{Ker}{\bf L}_{\alpha}^{\varepsilon}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$. Our next result shows that the moduli space $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is finite dimensional. Notice that both $\mathbf{P}$ and ${\bf L}_{\alpha}^{\varepsilon}$ are multi-degree differential operators, and hence we shall use the use the generalized notion of ellipticity provided by Douglis and Nirenberg [DN]. For the general theory of linear multi-degree elliptic differential operators we refer to [LM].
@@
$\mathbf{P}_{11}(f)$ $=-L_{\eta_{f}}J,$ $\mathbf{P}_{12}$ $=0,$ $\mathbf{P}_{21}(f)$ $=-i_{\eta_{f}}F_{A}$
@@
$o(\mathbf{P}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{P}_{12})$ $=0\leqslant 2-0,$ $o(\mathbf{P}_{21})$ $=1\leqslant 2-1$
@@
$o(\mathbf{L}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{12})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{13})$ $=1\leqslant 2-0$
@@
$u_{4}\omega=v\wedge a,\qquad\textrm{ for }a\in\Omega^{1}(X,E_{K}(\mathfrak{k}).$
@@
To finish, assume that $\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)(\dot{J},a,\dot{\psi})=0$. Then, $\sigma_{\mathbf{L}_{4}}(v)=0$ implies that there exists $x\in\Omega^{0}(X,E_{K}(\mathfrak{k})$ such that
@@
and therefore $\sigma_{\mathbf{L}_{1}}(v)=0$ implies $v\wedge v(\dot{J})=0$. By [F], there exists a smooth function $f\in C^{\infty}_{0}(\Sigma,\mathbb{R})$ such that
$\dot{J}=-fJv\otimes\omega^{-1}v$
@@
Finally, writting $\omega=\lambda v\wedge Jv$ and
$F_{A}=tv\wedge Jv\textrm{ for }t\in\Omega^{0}(X,E_{K}(\mathfrak{k}),$
@@
$(\dot{J},a,\dot{\psi})=\sigma_{{\bf P}}(v)(f,-x-\lambda^{-1}ft).$
Our strategy to build a complex structure induced by (3.7) on the moduli space is to work orthogonally to the image of the infinitesimal action operator $\mathbf{P}$ in (4.7) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.18). The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the two-form $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.17) is well defined on the cohomology $H^{1}(\mathcal{S}^{*})$ by Proposition 3.9. Consider the $L^{2}$-pairing on $\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(X,E_{K}(\mathfrak{k})$, the domain of the operator $\mathbf{P}$, induced by $\omega$ and $B$:
@@
$\widetilde{\mathbf{P}}^{*}=\delta\mu\circ\mathbb{J}\colon\mathcal{S}^{1}\to\mathcal{S}^{0}.$
@@
$\begin{split}\langle\widetilde{\mathbf{P}}^{*}v,(f,u)\rangle=\langle d\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(\mathbb{J}v),\zeta\rangle=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(\zeta\cdot(J,A,\psi),\mathbb{J}v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(\widetilde{\mathbf{P}}(f,u),v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,\widetilde{\mathbf{P}}(f,u)).\end{split}$
@@
[…]
The key condition on the solution $(J,A,\psi)$ of (4.3) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}$. Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\operatorname{ker}\mathcal{L}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$. We build on the following technical result, which shows in particular that the subspace of null vectors in $\operatorname{Im}\mathbf{P}$ is finite-dimensional. Using the $L^{2}$ norm (4.8), we extend the domain of $\mathcal{L}$ to an appropriate Sobolev completion.
@@
The operator $\mathcal{L}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\operatorname{ker}\mathcal{L}$ are smooth.
@@
$o(\widetilde{\mathcal{L}}_{11})$ $=4\leqslant 4-0,$ $o(\widetilde{\mathcal{L}}_{12})$ $=2\leqslant 4-0,$
$o(\widetilde{\mathcal{L}}_{21})$ $=2\leqslant 4-2,$ $o(\widetilde{\mathcal{L}}_{22})$ $=2\leqslant 4-2.$
@@
which is clearly invertible. Consequently, by the general theory of linear multi-degree elliptic differential operators (see [LM]) it follows that $\widetilde{\mathcal{L}}$ is Fredholm and elements in $\operatorname{ker}\widetilde{\mathcal{L}}$ are smooth. The result follows now from $\operatorname{ker}\mathcal{L}=\mathbf{T}^{-1}\operatorname{ker}\widetilde{\mathcal{L}}$ and the fact that $\mathcal{L}$ is self-adjoint. ∎
Assuming that $\operatorname{ker}\mathcal{L}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$-orthogonal decomposition

### v2

@@ 3.3. Minimal coupling and the Hamiltonian action
For the second equality, we use the identity $L_{y}J=2i\bar{\partial}y^{1,0}-2i\partial y^{0,1}$ for any vector field $y$ on $\Sigma$, and therefore
$B(J\psi\wedge\psi(L_{y}J))$ $=2B(\psi^{1,0}\wedge\psi(\bar{\partial}y^{1,0}))+\textnormal{c.c.}$
$=2B(\psi^{1,0}\wedge(i_{y^{1,0}}(\bar{\partial}_{A}\psi^{1,0})+\bar{\partial}_{A}(i_{y^{1,0}}\psi)))+\textnormal{c.c.}$
$=-2\bar{\partial}(B(\psi^{1,0},\psi(y^{1,0})))+4B(\bar{\partial}_{A}\psi^{1,0},\psi(y^{1,0}))+\textnormal{c.c.},$
where as usual “c.c.” stands for complex conjugate. ∎
The $\widetilde{{\curly K}}-$action produces a ‘coupling term’ on the base $\Sigma$ (cf. Remark 4.2), which interacts with the scalar curvature of the metric $g=\omega(,J)$. In particular, as in [AGG1], the zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}$ correspond to the solutions of the coupled system of equations
(3.23) $\begin{split}d_{A}^{*}\psi&=0,\\ \varepsilon S_{g}-\alpha*d\left(B\left(\Lambda_{\omega}F_{A},*\psi\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
where $\chi(\Sigma)$ is the Euler characteristic of $\Sigma$ and $V=\int_{\Sigma}\omega$ is the total volume. In the next section we combine these equations with the flatness condition $F_{D}=0$ for the complex connection to introduce the universal Hitchin moduli space.
@@ 4.1. Universal moduli space of flat $\boldsymbol{G}$-connections
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Let $\curly D$} be the space of connections on $E_{G}$, equipped with the constant complex structure $\mathbf{J}$ given by (2.1). Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the Kähler fibration $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ defined in Section 3.2. Our aim in this section is to investigate the existence of a Kähler structure on $\curly{X}$ and on its quotients, by combining the minimal coupling associated with the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6, as discussed in Section 3.3, with symplectic reduction. Difficulties arise from Corollary 3.7, which shows that the associated symmetric tensor $\mathbf{g}_{\mathbb{J}}:=\sigma(,\mathbb{J})$ is negative semi-definite along the horizontal directions.
@@
In other words, the extension (4.1) is given by restriction of the natural map from holomorphic automorphisms of $\curly{X}$ which preserve the fibration structure, to holomorphic automorphisms of the base $\curly{J}$. It is important to observe that even though $\operatorname{Diff}_{0}(\Sigma)$ acts by holomorphic automorphisms of the base $\curly{J}$, this is a real infinite-dimensional Lie group, which does not admit a complexification (see e.g. [Do3]).
@@
As in the previous section, $\Omega_{\mathbb{J}}$ induces a structure of holomorphic symplectic fibration on $\curly{X}\to\curly{J}$ and a holomorphic Ehresmann connection, which in this case is simply the trivial connection. We will return to this structure when studying the universal Higgs bundle moduli space in Section 5. The following result is an immediate consequence of the proof of [AGG1, Proposition 1.6], but we sketch the proof in the present setup for the convenience of the reader.
@@
where $D\zeta\in\Omega^{0}(\Sigma,E_{G}(\mathfrak{g}))$ denotes the vertical part of the $G$-equivariant vector field $\zeta$ on the total space of $E_{G}$, with respect to the connection $D$.
@@
The result now follows because the infinitesimal action of $\zeta$ on $D$ is given by
$\zeta\cdot D=-d_{D}(D\zeta)-i_{p\zeta}F_{D}.\qed$
It is remarkable that, unlike in the moment-map calculation in Proposition 3.9, in the present setup the extension of the complex gauge group by diffeomorphisms does not produce any additional ‘coupling term’ in the base $\Sigma$ (cf. Remark 4.2). One can interpret this fact as a functorial property of flat connections under the action of diffeomorphisms on a surface. This allows us to define the universal moduli space of flat $G$-connections, as follows, and to impose the flatness condition for $D$ as an integrability condition in the next sections.
@@
$\mu_{\widetilde{\curly{G}}|\curly{X}^{*}}\colon\curly{X}^{*}\longrightarrow\operatorname{Lie}\widetilde{\curly{G}}^{*}$.
@@
By construction, $\mathcal{U}^{\operatorname{Flat}}(G)$ fibres over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
(4.2) $\mathcal{U}^{\operatorname{Flat}}(G)\longrightarrow\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
@@
For a $G$-bundle on a higher dimensional symplectic manifold $(M^{2n},\omega)$, consider the subgroup $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$ given by the preimage by $p$ of the group of Hamiltonian symplectomorphisms on $M$. Then the proof of [AGG1, Proposition 1.6] shows that the induced $\widetilde{{\curly K}}-$action on $\curly D$} is Hamiltonian, with equivariant moment map
@@
where $f\in C^{\infty}(M,\mathbb{R})$ is the unique smooth function satisfying $i_{y}\omega=df$ and $\int_{M}f\omega^{n}=0$, with $p(\zeta)=y$. In this case, there is a ‘coupling term’ on the base corresponding to the function
@@ 4.2. Universal moduli space of the coupled harmonic equations
In this section we combine the results of the previous two sections in order to define universal moduli spaces of solutions of the harmonicity equations (2.11), varying over Teichmüller space. As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
@@
A solution of the coupled harmonic equations with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$ is a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ that satisfies the equations
@@
As in Section 4.1, we consider the complex subspace ${\curly D}^{*}\subset{\curly D}$ of reductive flat connections. Then $\curly{X}^{*}:=\curly{J}\times{\curly D}^{*}$ is formally a complex submanifold of $\curly{X}$ preserved by the $\widetilde{{\curly K}}-$action (see Section 3.3), and inherits a minimal coupling structure $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ and moment map $\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}$. We define the universal Hitchin moduli space as the symplectic quotient
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}:=\mu_{\widetilde{{\curly K}}|\curly{X}^{*}}^{-1}(0)/\widetilde{{\curly K}}$.
By construction, $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ parametrizes solutions of the coupled harmonic equations (4.3) with coupling constant $\alpha>0$ and parameter $\varepsilon\in\{-1,1\}$, modulo the $\widetilde{{\curly K}}-$action, and inherits a natural presymplectic structure on its smooth locus. This structure is furthermore symplectic in the case $\varepsilon=-1$, by Lemma 3.8. There are natural maps (cf. (5.3))
$\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}\longrightarrow\mathcal{U}^{\operatorname{Flat}}(G)\longrightarrow\mathcal{T}.$
@@
(4.4) $(\dot{J},a,\dot{\psi})\in\mathcal{S}^{1}:=T_{J}\curly{J}\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k})).$
@@
We denote by ${\bf L}_{\alpha}^{\varepsilon}(\dot{J},a,\dot{\psi})$ the differential operator defined by the left-hand side of equations (4.5). We turn next to the study of the infinitesimal action, in order to define a complex of differential operators. From the proof of Proposition 3.9, using the connection $A$ we can identify elements $\zeta\in\operatorname{Lie}\widetilde{{\curly K}}$ with pairs
$\zeta\cong(f,u)\in\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(\Sigma,E_{K}(\mathfrak{k})),$
@@
$\mathcal{S}^{2}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus(\Omega^{2}(\Sigma,E_{K}(\mathfrak{k})))^{\oplus 3},$
@@
[…]
The cohomology $H^{1}(\mathcal{S}^{*}):=\frac{\ker{\bf L}_{\alpha}^{\varepsilon}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$. Our next result shows that the moduli space $\mathcal{U}^{\operatorname{Harm}}(G)_{\alpha}^{\varepsilon}$ is finite dimensional. Notice that both $\mathbf{P}$ and ${\bf L}_{\alpha}^{\varepsilon}$ are multi-degree differential operators, and hence we shall use the generalized notion of ellipticity provided by Douglis and Nirenberg [DN]. For the general theory of linear multi-degree elliptic differential operators we refer to [LM].
@@
$\mathbf{P}_{11}(f)$ $=-L_{\eta_{f}}J,$ $\mathbf{P}_{12}$ $=0,$ $\mathbf{P}_{21}(f)$ $=-i_{\eta_{f}}F_{A},$
@@
$o(\mathbf{P}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{P}_{12})$ $=0\leqslant 2-0,$ $o(\mathbf{P}_{21})$ $=1\leqslant 2-1,$
@@
$o(\mathbf{L}_{11})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{12})$ $=2\leqslant 2-0,$ $o(\mathbf{L}_{13})$ $=1\leqslant 2-0,$
@@
$u_{4}\omega=v\wedge a,\qquad\textrm{ for }a\in\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}).$
@@
To finish, assume that $\sigma_{{\bf L}_{\alpha}^{\varepsilon}}(v)(\dot{J},a,\dot{\psi})=0$. Then, $\sigma_{\mathbf{L}_{4}}(v)=0$ implies that there exists $x\in\Omega^{0}(\Sigma,E_{K}(\mathfrak{k})$ such that
@@
and therefore $\sigma_{\mathbf{L}_{1}}(v)=0$ implies $v\wedge v(\dot{J})=0$. By [Fu], there exists a smooth function $f\in C^{\infty}_{0}(\Sigma,\mathbb{R})$ such that
$\dot{J}=-fJv\otimes\omega^{-1}v.$
@@
Finally, writing $\omega=\lambda v\wedge Jv$ and
$F_{A}=tv\wedge Jv\textrm{ for }t\in\Omega^{0}(\Sigma,E_{K}(\mathfrak{k}),$
@@
$(\dot{J},a,\dot{\psi})=\sigma_{{\bf P}}(v)(f,-x-\lambda^{-1}ft).\qed$
Our strategy to build a complex structure induced by (3.7) on the moduli space is to work orthogonally to the image of the infinitesimal action operator $\mathbf{P}$ in (4.7) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.18). The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the 2-form $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ in (3.17) is well defined on the cohomology $H^{1}(\mathcal{S}^{*})$ by Proposition 3.9. Consider the $L^{2}$-pairing on $\mathcal{S}^{0}:=C_{0}^{\infty}(\Sigma,\mathbb{R})\oplus\Omega^{0}(\Sigma,E_{K}(\mathfrak{k})$, the domain of the operator $\mathbf{P}$, induced by $\omega$ and $B$:
@@
$\widetilde{\mathbf{P}}^{*}=\delta\mu\circ\mathbb{J}\colon\mathcal{S}^{1}\longrightarrow\mathcal{S}^{0}.$
@@
$\langle\widetilde{\mathbf{P}}^{*}v,(f,u)\rangle$ $=\langle d\mu_{\widetilde{{\curly K}}}^{\mathbb{J}}(\mathbb{J}v),\zeta\rangle=\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(\zeta\cdot(J,A,\psi),\mathbb{J}v)$
$=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(\widetilde{\mathbf{P}}(f,u),v)=\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,\widetilde{\mathbf{P}}(f,u)).\qed$
@@
(4.11) $\begin{array}[]{cccl}\mathcal{L}\colon&\mathcal{S}^{0}&\longrightarrow&\mathcal{S}^{0}\\ &(f,u)&\longmapsto&\widetilde{\mathbf{P}}^{*}\circ\widetilde{\mathbf{P}}(f,u).\end{array}$
The key condition on the solution $(J,A,\psi)$ of (4.3) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}$. Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\ker\mathcal{L}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$. We build on the following technical result, which shows in particular that the subspace of null vectors in $\operatorname{Im}\mathbf{P}$ is finite-dimensional. Using the $L^{2}$ norm (4.8), we extend the domain of $\mathcal{L}$ to an appropriate Sobolev completion.
@@
The operator $\mathcal{L}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\ker\mathcal{L}$ are smooth.
@@
$o(\widetilde{\mathcal{L}}_{11})$ $=4\leqslant 4-0,\quad$ $\quad o(\widetilde{\mathcal{L}}_{12})$ $=2\leqslant 4-0,$
$o(\widetilde{\mathcal{L}}_{21})$ $=2\leqslant 4-2,$ $o(\widetilde{\mathcal{L}}_{22})$ $=2\leqslant 4-2,$
@@
which is clearly invertible. Consequently, by the general theory of linear multi-degree elliptic differential operators (see [LM]) it follows that $\widetilde{\mathcal{L}}$ is Fredholm and elements in $\ker\widetilde{\mathcal{L}}$ are smooth. The result follows now from $\ker\mathcal{L}=\mathbf{T}^{-1}\ker\widetilde{\mathcal{L}}$ and the fact that $\mathcal{L}$ is self-adjoint. ∎
Assuming that $\ker\mathcal{L}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$-orthogonal decomposition